OpenStudy (elisaneedshelp):

Please if someone would be nice enough to help because im failing geometry and im trying to bring my grade up and i only have today! A window frame that seems rectangular has height 408 cm, length 306 cm, and one diagonal with length 525 cm. Is the window frame really rectangular? Explain.

OpenStudy (anonymous):

A is Length so A^2 would be 306^2; B is Height so B^2 would be 408^2 The Pythagorean Theorem is A^2 + B^2 = C^2

OpenStudy (anonymous):

it would be the idea that rectangles have right angles so you would use the pythagorean theorem to test this a^2 + b^2 = c^2 c = sqrt ( 408^2 + 306^2 ) = ? would it equal 525 ?

OpenStudy (elisaneedshelp):

why square root?

OpenStudy (anonymous):

because c^2 = a^2 + b^2 right? :) then to get c alone from c^2 you would sqrt both sides so c = sqrt ( a^2 + b^2 )

OpenStudy (elisaneedshelp):

can you give me an example :) lol sorry :/

OpenStudy (anonymous):

it's alright :) like if you wanted to check if sides 3 and 4 made a rectangle with diagonal 6 then a^2 + b^2 = c^2 c = sqrt (a^2 + b^2) = sqrt ( 3^2 + 4^2 ) = ? would it equal 6 ? if it does, it is a rectangle. if it does not, then it is not a rectangle

OpenStudy (elisaneedshelp):

Nope it equals 5 so then its not a rectangle

OpenStudy (anonymous):

yes :) so then you would use the same idea for your own question :)

OpenStudy (elisaneedshelp):

thank you :) can you help me with this one , I have to find Y

OpenStudy (elisaneedshelp):

Omg thanks my first testimonial :'D

OpenStudy (anonymous):

pythagorean theorem c would = 25 a would = -7 so 25^2 = (-7)^2 + y^2 solve for y

OpenStudy (elisaneedshelp):

\[-7^{2}+b ^{2}=25^{2}\] \[-49+b ^{2}=625\] +49 +49 b= 674

OpenStudy (anonymous):

it was 7^2 + b^2 = 25^2 so 49 + b^2 = 625 -49 -49 b^2 = 576 so b = sqrt 576 = ?

OpenStudy (anonymous):

ah right, I saw what you did when the negative is in the parenthesis, it is included in the squaring so (-7)^2 = 49 which is different from - (7)^2 = -49 we wanted (-7)^2 = 49

OpenStudy (elisaneedshelp):

But when i put -7^2 it gave me -49

OpenStudy (anonymous):

(-7) times (-7) = 49 because a negative times a negative = positive

OpenStudy (elisaneedshelp):

Ohhh..........i didnt know it had to be in ( )

OpenStudy (anonymous):

@Emz.Yaoi i just want to say. You should considering becoming a teacher lol

OpenStudy (elisaneedshelp):

Lol ikr

OpenStudy (anonymous):

Thanks @BeautifulSoulღ :) I'll be an ambassador of this website soon :P that'll be good enough for now

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

@ElisaNeedsHelp so do you understand how to do your problem now? :)

OpenStudy (elisaneedshelp):

Yes :) i was just copying down the other problem that you had explained to me

OpenStudy (anonymous):

coolios ^_^